OpenStudy (anonymous):

If g equals 3, what is the value of 2 g plus fifty-three? fifty-nine one hundred six one hundred fifty-nine three hundred eighteen

OpenStudy (anonymous):

we have, g=3 now, 2g = 2*3 =6 also, 2g+53 = 6+53=59